Complete the table:

Compound Mass Moles Number of molecules
CO2CO2 _____ 0.0193 _____
COCO _____ 0.0104 _____
BrIBrI 31.0 mgmg _____ _____
CF2Cl2CF2Cl2 1.44 kgkg _____ _____
Answer #1

1.

CO2CO2

we have moles for  CO2CO2 = 0.0193

so we know  

no of moles = mass of compound / molar mass of compound

0.0193 = mass / 88 (molar mass of CO2CO2 = 12+16+16+12+16+16 = 88 gm/mol)

mass = 1.6984 gm

we know 1 mole of any compound contain 6.02 x 1023 molecules

we have mole = 0.0193

so no of moleculs = no of moles x 6.02 x 1023

no of moleculs = 0.0193 x 6.02 x 1023 = 0.116 x 1023

2.

COCO

we have moles for  COCO= 0.0104

so we know  

no of moles = mass of compound / molar mass of compound

0.0104 = mass / 56 (molar mass of COCO = 12+16+12+16= 56 gm/mol)

mass = 0.5824  gm

we know 1 mole of any compound contain 6.02 x 1023 molecules

we have mole = 0.0104

so no of moleculs = no of moles x 6.02 x 1023

no of moleculs = 0.0104 x 6.02 x 1023 = 0.062 x 1023

3.

BrIBrI

we have mass of BrIBrI = 31 mg = 31 x 10-3 gm (1 mg = 10-3 g)

so we know  

no of moles = mass of compound / molar mass of compound

no of moles = 31 x 10-3 / 414 (molar mass of BrIBrI = 80+ 127+80+127= 414 gm/mol)

no of moles = 0.074 x 10-3

we know 1 mole of any compound contain 6.02 x 1023 molecules

we have mole =0.074 x 10-3

so no of moleculs = no of moles x 6.02 x 1023

no of moleculs = 0.074 x 10-3x 6.02 x 1023 = 0.445 x 1020

4.

CF2Cl2CF2Cl2

we have mass of CF2Cl2CF2Cl2 = 1.44 Kg = 1.44 x 103 gm (1 Kg = 103 g)

so we know  

no of moles = mass of compound / molar mass of compound

no of moles = 1.44 x 103  / 242 (molar mass of CF2Cl2CF2Cl2 = 242 gm/mol)

no of moles = 5.95

we know 1 mole of any compound contain 6.02 x 1023 molecules

we have mole =5.95

so no of moleculs = no of moles x 6.02 x 1023

no of moleculs = 5.95x 6.02 x 1023 = 35.819 x 1023